Grade 12Magical Mathematics[Interesting Approach]

Find all integers `n` for which (n^2+13n+2)/(3n+5) also an Integer ?

And Please give me some hint that How to slove such type of Questions ?

Surya Anuraag Duvvuri
15 Years agoGrade 12

1 Answer

SAGAR SINGH - IIT DELHI
15 Years ago

Dear student,

Here goes the hint:

6n² + 13n + 6 = 0
n² + 13/12n = - 1 + (13/12)²
n² + 13/12n = - 144/144 + 169/144
(n - 13/12)² = 25/144
n - 13/12 = ± 5/12

n = 5/12 + 13/12, n = 18/12, n = 3/2
n = - 5/12 + 13/12, n = 8/12, n = 2/3

Answer: n = 3/2, 2/3

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi